# Relaxation time approximation revisited and pole/cut structure in retarded correlators

### v1

@@ V The novel relaxation time approximation
note that the weight function is the same as the one in Eq.(9), as it should be. In Eq.(42), the novel RTA is constructed by adding counter terms to the RTA, but it can also be derived through an alternative method, as shall be given below. To proceed, we expand $\chi(p)$ as the linear combination of the eigenfunctions of $\mathcal{L}_{0}$ as
@@
where the second term on the right-hand side precisely constitutes the counter term required to restore collision invariance. Hence, Eq.(V) encompasses Eq.(42) as a particular case. Compared to the constructions in [3, 26, 35], the method introduced here embodies the essence of truncation that leads to RTA and is capable of accommodating a broader spectrum of modified RTAs. By increasing $N$, we can incorporate additional eigenvalues and eigenfunctions into the original RTA if available. This approach is beneficial when our knowledge is limited to a finite set of eigenvalues and eigenfunctions (these can be determined numerically at least [6]), when we attempt to integrate the relevant information into the RTA. Furthermore, the model in Eq.(V), serving as an intermediate between the RTA and the full linearized collision operator, can be fine-tuned to balance accuracy and simplicity, allowing for a judicious compromise.
@@ VI Summary and outlook
In this paper, we revisit the widely-used relaxation time approximation within the linearized Boltzmann equation. According to the mathematical analysis on the eigenspectrum of linearized Boltzmann collision operator $\mathcal{L}_{0}$, the RTA model is justified only for hard interactions, thereby ruling out the energy-dependent parametrization in Eq.(11). The consideration is grounded on mathematical aspects, and Eq.(11) can effectively serve as a convenient parameterized model. Furthermore, Eq.(11) sheds light on the redefinition of linearized collision operator denoted as $\mathcal{L}_{\alpha}$. We demonstrate that, using only physical arguments, one can related the eigenspectrum properties of the redefined $\mathcal{L}_{\alpha}$ to the softness of interactions through the available mathematical analysis on $\mathcal{L}_{0}$. This, in our view, is nontrivial, as intuitive physical considerations can yield meaningful mathematical insights without the need for sophisticated analysis. Thus far, we have completed the summary for Q1 and Q3.
Regarding Q2, we find that focusing on $\mathcal{L}_{0}$ and its eigenspectrum properties is the simplest approach. When interactions are hard, the RTA is well-defined, and the analytical properties of retarded correlators, as detailed in [1], are applicable. There is a gap between the branch-cut lines and the real axis $\text{Im}\,\omega=0$. Therefore, the gapless hydrodynamic modes are well-defined low energy degrees of freedom when $k$ is small. However, according to mathematical derivations, relativistic interactions are often soft, leading us to focus on the alternative scenario in most cases. In scenarios with soft interactions, the RTA is no longer well-justified, and the branch-cut lines form a strip that touches the origin. Hence, the dominating non-analytical structure turns into the branch-cut or the non-hydrodynamic modes. Our conclusion is consistent with the previous related studies with the comparison details elaborated in the main text. Note that if the theory is not conformal, or particles constituting the system are massive, the non-analytical structures would be more complicated and richer, as detailedly exhibited in Sec.IV. We also provide a derivation of the novel RTA to restores collision invariance by truncating the linearized collision operator, rather than by adding counter terms.

### v2

@@ V The novel relaxation time approximation
note that the weight function is the same as the one in Eq.(II), as it should be. In Eq.(37), the novel RTA is constructed by adding counter terms to the RTA, but it can also be derived through an alternative method, as shall be given below. To proceed, we expand $\chi(p)$ as the linear combination of the eigenfunctions of $\mathcal{L}_{0}$ as
@@
where the second term on the right-hand side precisely constitutes the counter term required to restore collision invariance. Hence, Eq.(V) encompasses Eq.(37) as a particular case. One can also show that the collision invariance is restored for $N>6$. Compared to the constructions in [27, 34, 37], the method introduced here embodies the essence of truncation that leads to RTA and is capable of accommodating a broader spectrum of modified RTAs. By increasing $N$, we can incorporate additional eigenvalues and eigenfunctions into the original RTA if available. This approach is beneficial when our knowledge is limited to a finite set of eigenvalues and eigenfunctions (these can be determined numerically at least [13]), when we attempt to integrate the relevant information into the RTA. Furthermore, the model in Eq.(V), serving as an intermediate between the RTA and the full linearized collision operator, can be fine-tuned to balance accuracy and simplicity, allowing for a judicious compromise.
@@ VI Summary and outlook
In this paper, we revisit the widely-used relaxation time approximation within the linearized Boltzmann equation. According to the mathematical analysis on the eigenspectrum of linearized Boltzmann collision operator $\mathcal{L}_{0}$, the RTA model is justified only for hard interactions, thereby ruling out the energy-dependent parametrization in Eq.(8). The consideration is grounded on mathematical aspects, and Eq.(8) can effectively serve as a convenient parameterized model. Furthermore, Eq.(8) sheds light on the redefinition of linearized collision operator denoted as $\mathcal{L}_{\alpha}$. We demonstrate that, using only physical arguments, one can related the eigenspectrum properties of the redefined $\mathcal{L}_{\alpha}$ to the softness of interactions through the available mathematical analysis on $\mathcal{L}_{0}$. This, in our view, is nontrivial, as intuitive physical considerations can yield meaningful mathematical insights without the need for sophisticated analysis. Thus far, we have completed the summary of Q1. We also provide a derivation of the novel RTA to restores collision invariance by truncating the linearized collision operator, rather than by adding counter terms.
Regarding Q2, we find that focusing on $\mathcal{L}_{0}$ and its eigenspectrum properties is the simplest approach. When interactions are hard, the RTA is well-defined, and the analytical properties of retarded correlators, as detailed in [7], are applicable. There is a gap between the branch-cut lines and the real axis $\text{Im}\,\omega=0$. Therefore, the gapless hydrodynamic modes are well-defined low energy degrees of freedom when $k$ is small. However, according to mathematical derivations, relativistic interactions are often soft, leading us to focus on the alternative scenario in most cases. In scenarios with soft interactions, the RTA is no longer well-justified, and the dominating long-lived non-analytical structure turns into the branch-cut or the non-hydrodynamic modes. Our conclusion is consistent with the previous related studies with the comparison details elaborated in the main text. Note that if particles constituting the system are massive or the perturbations are inhomogeneous, the non-analytical structures would be more complicated and richer.